# Audio xLSTMs: Learning Self-supervised audio representations with xLSTMs

### v1

@@ Abstract
While the transformer has emerged as the eminent neural architecture, several independent lines of research have emerged to address its limitations. Recurrent neural approaches have also observed a lot of renewed interest, including the extended long short-term memory (xLSTM) architecture, which reinvigorates the original LSTM architecture. However, while xLSTMs have shown competitive performance compared to the transformer, their viability for learning self-supervised general-purpose audio representations has not yet been evaluated. This work proposes Audio xLSTM (AxLSTM), an approach to learn audio representations from masked spectrogram patches in a self-supervised setting. Pretrained on the AudioSet dataset, the proposed AxLSTM models outperform comparable self-supervised audio spectrogram transformer (SSAST) baselines by up to 20% in relative performance across a set of ten diverse downstream tasks while having up to 45% fewer parameters.
Index Terms:
xLSTM, self-supervised learning, audio representation learning.
I Introduction
@@
In lieu of their excellent generalisation capabilities and domain and data agnostic nature, transformers [1] and their successors have seen widespread adoption. Furthermore, transformers, together with masked predictive modelling, have emerged as a key driving force behind several prominent advancements in the realm of unsupervised and self-supervised representation learning for NLP [2], computer vision [3, 4] and audio and speech processing [5, 6, 7, 8, 9]. However, the race to find alternatives to scaled dot-product attention at the heart of transformers has heated up, with a lot of emphasis on finding sub-quadratic appromixations of the attention operation [10] as well as token mixing [11].
@@
This search for alternatives has led us back to recurrent neural approaches. Before the advent of transformers, sequence modeling was predominantly done using recurrent neural networks (RNNs), such the long short-term memory (LSTM) [12]. Recurrent models offer several advantages over transformers: they scale linearly with respect to sequence length and they have lower runtime memory requirements since storing the entire key-value (KV) cache is not necessary. The search has led us to new approaches such as state space models (SSMs) [13, 14, 15, 16, 17], which are a family of sequence models that lie at the intersection of convolutional neural networks, RNNs and classical state spaces. Several variants of SSMs have since been proposed, showing competitive performance and scalability versus transformers in several domains, including long sequence modelling [16, 17], computer vision [18] as well as audio [19].
@@
Compared to the transformer architecture, LSTMs suffer from several key drawbacks: (i) inability to revise storage decisions, (ii) compressing information into a scalar cell state, which impacts performance on rare input tokens, and (iii) memory mixing and the resulting lack of parallelizability. Recently, [20] proposed the extended long-short term memory (xLSTM) neural architecture, which revitalises the original LSTM architecture, leveraging the latest techniques and tricks learned from years of transformer and large language modelling research. xLSTM incorporates exponential gating, improved normalization and stabilization techniques while removing traditional memory mixing, resulting in two new fundamental building blocks: sLSTM and mLSTM. xLSTMs have demonstrated performance on-par with or better than transformers and Mamba for large language modelling [20] as well as a generic vision backbone trained on non-overlapping image patches [21], while also demonstrating superior sequence length extrapolation capabilities. However, the ability of xLSTMs to learn general-purpose audio representations in a self-supervised setting is yet to be thoroughly evaluated.
@@
In this work, we propose Audio xLSTMs (AxLSTMs) for learning self-supervised general-purpose audio representations within a masked modelling framework from spectrograms patches. Pretrained on AudioSet [22], AxLSTMs consistently outperform comparable self-supervised audio spectrogram transformer (SSAST) [7] based baselines on ten varied downstream tasks, improving aggregate performance by over 20% while having over 45% fewer parameters. Code and pretrained models will be released publicly.
II Method
II-A Prerequisites: xLSTM and the mLSTM block
As previously discussed, xLSTM [20] proposes two new building blocks, sLSTM and mLSTM. In line with Vision-LSTM [21], the parallelizable mLSTM module is the fundamental building block of our approach, and thus is the focus of our discussion. The mLSTM block addresses the issues with the original LSTM cell by utilizing a matrix memory cell $C\in\mathbb{R}^{d\times d}$. mLSTM stores a key-value vector pair $k_{t},v_{t}\in\mathbb{R}^{d}$ at timestep $t$, and the relevant value vector is retrieved later using a query $q_{t+\tau}\in\mathbb{R}^{d}$. The forward pass for mLSTM is as follows:
@@ II-A Prerequisites: xLSTM and the mLSTM block
where $C_{t}$ is the matrix memory cell, $n_{t}$ and $h_{t}$ represent the normalizer state and the hidden state, and $i_{t},f_{t}$ and $o_{t}$ represent the input, forget and the output gates, respectively. Weights $W_{q},W_{k}$ and $W_{v}$ are learnable projection matrices for vectors query $q$, key $k$ and value $v$, respectively. Since there is no memory mixing in mLSTMs, multiple memory cells and multiple heads are equivalent, and the forward pass can be parallelized. Further, mLSTM uses exponential gating to mitigate the inability of LSTMs to revise storage decisions.
II-B AxLSTM: Modelling masked patches with xLSTMs
Creating and masking patches: We compute non-overlapping patches of shape $t\times f$ from input spectrogram $\mathbf{x}\in\mathbb{R}^{\mathtt{t}\times\mathtt{f}}$, yielding $\mathbf{x}_{p}\in\mathbb{R}^{N\times(t\cdot f)}$ patches. After flattening these patches and projecting them linearly to a $\mathbb{R}^{N\times d_{m}}$ dimensional space, we encode positional information by adding fixed sinusoidal positional. A representative class token is then added to the beginning of the sequence, similar to [23, 7]. We then proceed to randomly mask 50% of the input patches using an unstructured masking strategy, and replace these masked patches with a learnable mask token. Thus, input to the encoder is:
@@ II-B AxLSTM: Modelling masked patches with xLSTMs
Encoding and Reconstruction: These partially masked patches are now fed to the encoder, yielding encoded representations $\mathbf{z}=\mathtt{enc}(\mathbf{x}^{\prime}),\mathbf{z}\in\mathbb{R}^{(N+1)\times d_{m}}$. The encoder is a stack of mLSTM blocks [20, 21] (Fig. 1), and each block expands the $d_{m}$ dimensional input by an expansion factor $E=2$ before projecting it back to $d_{m}$. This yields blocks that have fewer parameters than a corresponding transformer block. After encoding, a single hidden layer MLP is used to reconstruct patches from encoded representation $\mathbf{z}$. :
@@
where $\mathtt{Linear}_{d}$ is a parameterized linear projection to dimensions $d$, and $\sigma$ denotes the GELU non-linear activation function. The reconstructed output $y$ is obtained after removing the $\mathtt{cls}$ token from $\mathbf{y^{\prime}}$. The mean square error between the original input patches and the reconstructions is used for pretraining. Further details can be found in Section III-B.
III Experiments
TABLE I: Comparing AxLSTMs with popular self-supervised audio representations. We used pretrained models from cited papers to extract fixed feature vectors and conducted our own downstream experiments. Directly comparable SSAST and AxLSTM models are highlighted with similar colours. LS, AS, VP, LL stand for LibriSpeech, AudioSet, VoxPopuli and LibriLight datasets, respectively. SSAST [7] was trained on AS+LS. *includes decoder parameters
Model Data # Params BO CD ESC-50 LC Mri-S Mri-T NS-5h SC-5h F50K VL $s(m)$
Naive Baselines
HEAR-Naive [24] - - 52.6±2.4 30.9±0.8 5.8±0.2 33.5±1.1 38.0±1.3 36.4±1.9 18.6±4.4 8.5±0.4 7.1±0.2 11.2±0.5 5.2±0.8
Supervised
PaSST-Base [25] AS 86 M 94.9±0.5 61.0±0.3 94.8±0.3 60.1±0.2 96.5±0.1 87.6±0.6 23.3±0.9 66.6±1.4 64.2±0.1 25.5±0.8 74.7±0.4
@@ III Experiments
W2V2-base [5] LS 94.4 M 74.0±1.0 46.4±0.3 31.1±0.4 51.2±0.2 77.3±0.2 55.1±0.3 7.4±0.8 90.8±0.3 18.1±0.1 35.5±0.8 44.1±0.2
W2V2-large [5] VP 315.4 M 93.1±0.7 66.9±0.4 60.1±0.5 62.4±0.3 93.9±0.1 77.4±0.2 42.0±1.0 87.6±0.5 34.2±0.1 53.6±1.0 75.1±0.4
WavLM-base [9] LS 94.4 M 89.4±0.7 56.3±0.2 46.6±0.4 63.2±0.3 95.1±0.1 83.4±0.2 37.3±0.8 57.2±0.8 29.9±0.1 22.6±0.6 61.7±0.2
WavLM-large [9] Mix 315.4 M 96.4±0.5 57.2±0.2 47.9±0.4 61.1±0.3 96.8±0.1 89.5±0.1 53.7±0.5 46.2±0.8 29.0±0.1 23.7±0.9 65.1±0.2
HuBERT-base [6] LS 94.4 M 92.1±0.6 70.8±0.2 57.8±0.6 56.5±0.3 94.4±0.1 84.9±0.3 19.4±0.7 93.2±0.1 32.3±0.1 61.8±0.6 74.5±0.2
HuBERT-large [6] LL 315.4 M 94.1±0.7 70.7±0.1 60.3±0.4 59.9±0.2 95.3±0.1 83.5±0.3 19.3±0.8 83.2±0.7 31.5±0.1 66.1±0.9 74.4±0.3
BEATs-iter3 [26] AS 90 M 94.0±0.8 67.3±0.2 83.7±0.3 68.0±0.2 94.7±0.1 95.8±0.1 69.4±0.8 85.2±0.3 53.6±0.2 38.5±1.0 87.6±0.3
AudioMAE [8] AS 86 M 93.7±0.6 68.2±0.2 60.6±0.4 42.2±0.2 89.2±0.2 86.6±0.2 64.5±0.8 28.6±1.5 37.9±0.1 29.7±1.0 64.3±0.3
MWMAE-Tiny [27] AS 12.6 M* 93.3±1.0 64.4±0.2 71.9±0.5 65.5±0.3 97.1±0.1 97.6±0.1 68.1±0.4 77.0±0.6 43.4±0.1 28.6±1.1 80.7±0.3
MWMAE-Base [27] AS 92.5 M* 96.0±0.5 73.1±0.3 81.2±0.4 68.8±0.2 97.4±0.1 97.9±0.1 69.3±0.6 90.9±0.2 51.2±0.2 44.2±0.9 91.3±0.2
SSAM-Tiny [19] AS 4.8 M 93.7±0.8 61.8±0.3 70.6±0.2 59.2±0.4 97.1±0.1 94.9±0.1 62.0±0.7 74.8±0.4 41.3±0.2 27.8±1.0 76.3±0.2
SSAM-Small [19] AS 17.9 M 94.0±0.7 67.5±0.2 78.7±0.6 60.5±0.3 97.5±0.1 96.7±0.1 66.3±0.8 83.7±0.3 48.5±0.1 39.6±0.7 84.4±0.3
SSAM-Base [19] AS 69.3 M 93.2±1.1 70.3±0.2 81.0±0.3 63.5±0.2 97.7±0.1 96.9±0.1 70.5±0.5 87.9±0.3 52.2±0.1 50.4±0.7 89.7±0.3
@@
SSAST [7] Mix 89 M 93.4±0.9 56.5±0.2 68.4±0.4 60.7±0.3 96.7±0.1 96.3±0.1 66.8±0.7 53.5±1.3 38.2±0.1 28.5±0.9 73.1±0.2
SSAST-Tiny AS 5.4 M 90.4±0.7 46.9±0.2 42.4±0.6 42.7±0.2 95.7±0.1 94.3±0.1 61.2±0.5 50.6±1.6 24.6±0.1 13.8±1.0 56.0±0.2
SSAST-Small AS 21.5 M 93.2±0.5 51.6±0.2 50.1±0.6 50.0±0.3 96.2±0.1 95.0±0.1 63.8±0.4 58.3±1.2 31.6±0.1 15.6±0.7 63.4±0.3
SSAST-Base AS 85.7 M 93.1±0.7 56.0±0.4 59.6±0.7 52.9±0.3 96.6±0.1 96.2±0.2 64.6±0.8 66.1±1.0 37.5±0.1 19.2±0.9 69.2±0.3
@@
AxLSTM-Tiny AS 2.9 M 93.1±0.7 53.7±0.4 58.3±0.3 54.2±0.4 96.4±0.1 94.2±0.1 61.6±0.7 62.5±1.1 35.1±0.1 20.4±1.0 67.0±0.3
AxLSTM-Small AS 11.2 M 94.2±0.4 63.4±0.2 72.2±0.5 61.7±0.4 97.2±0.0 96.4±0.1 71.7±0.4 76.0±0.9 43.6±0.1 31.3±0.5 79.8±0.2
AxLSTM-Base AS 44.0 M 93.9±0.6 66.1±0.2 75.9±0.5 63.0±0.3 97.0±0.1 96.9±0.1 72.1±0.6 80.4±0.3 47.1±0.1 35.8±0.8 83.1±0.2
III-A Datasets
Pretraining: For pretraining, we use the AudioSet dataset (AS) [22] which has roughly 2 million 10-second weakly labelled YouTube clips and over 5000 hours of audio data spanning 527 classes.
Downstream Evaluation: Following recent works [27, 19], we use a subset of the tasks proposed as a part of the HEAR benchmark [24] for downstream evaluation, comprising of the following ten diverse audio tasks: Beijing Opera, Crema-D, ESC-50, LibriCount, Mridangam Stroke and Tonic, NSynth Pitch 5h, Speech Commands 5h, FSD50K and VoxLingua107.
III-B Implementation details

### v3

# AxLSTMs: learning self-supervised audio representations with xLSTMs

@@ Abstract
While the transformer has emerged as the eminent neural architecture, several independent lines of research have emerged to address its limitations. Recurrent neural approaches have observed a lot of renewed interest, including the extended long short-term memory (xLSTM) architecture, which reinvigorates the original LSTM. However, while xLSTMs have shown competitive performance compared to the transformer, their viability for learning self-supervised general-purpose audio representations has not been evaluated. This work proposes Audio xLSTM (AxLSTM), an approach for learning audio representations from masked spectrogram patches in a self-supervised setting. Pretrained on the AudioSet dataset, the proposed AxLSTM models outperform comparable self-supervised audio spectrogram transformer (SSAST) baselines by up to 25% in relative performance across a set of ten diverse downstream tasks while having up to 45% fewer parameters.
keywords:
xLSTM, self-supervised learning, audio representation learning
1 Introduction
@@
In lieu of their excellent generalisation capabilities and domain and data agnostic nature, transformers [1] and their successors have seen widespread adoption. Furthermore, transformers, together with masked predictive modelling, have emerged as a key driving force behind several prominent advancements in the realm of unsupervised and self-supervised representation learning for NLP [2], computer vision [3] and audio and speech processing [4, 5, 6, 7, 8]. Recently, however, finding alternatives to scaled dot-product attention has garnered significant interest, with a lot of emphasis on finding sub-quadratic appromixations of the attention operation [9] as well as token mixing [10].
@@
Before the advent of transformers, sequence modeling was predominantly done using recurrent neural networks (RNNs). Recurrent models offer several advantages over transformers: they scale linearly with respect to sequence length and they have lower runtime memory requirements since storing the entire key-value (KV) cache is not necessary. State-space models (SSMs) [11], which are a family of sequence models that lie at the intersection of convolutional neural networks, RNNs and classical state spaces, are the most popular amongst recent recurrent approaches. Several variants of SSMs have been proposed, showing competitive performance and scalability versus transformers in several domains, including long sequence modelling [12], computer vision [13] as well as audio [14].
@@
However, prior to transformers and state-space models, LSTMs [15] were the go to neural architecture for sequence modeling. Compared to transformers, LSTMs suffer from several key drawbacks: (i) inability to revise storage decisions, (ii) compressing information into a scalar cell state, which impacts performance on rare input tokens, and (iii) memory mixing and the resulting lack of parallelizability. Recently, [16] proposed the extended long-short term memory (xLSTM) neural architecture, which revitalises the original LSTM architecture, leveraging the latest techniques and tricks learned from years of transformer and large language modelling research. xLSTM incorporates exponential gating, improved normalization and stabilization techniques while removing traditional memory mixing, resulting in two new fundamental building blocks: sLSTM and mLSTM. xLSTMs have demonstrated on-par or better performance than transformers for large language models [16] as well as patch-based image recognition [17], while also demonstrating superior sequence length extrapolation capabilities. However, the ability of xLSTMs to learn general-purpose audio representations in a self-supervised setting is yet to be thoroughly evaluated. In this work, we propose Audio xLSTMs (AxLSTMs) for learning self-supervised general-purpose audio representations within a masked modelling framework from spectrograms patches. Pretrained on AudioSet [18], AxLSTMs consistently outperform comparable self-supervised audio spectrogram transformer (SSAST) [6] based baselines on ten varied downstream tasks, even matching the performance of more recent and involved masked modeling approaches, such as BEATs [19]. Code and pretrained models can be found at https://github.com/SarthakYadav/axlstm-official.
@@
2 Method
2.1 Prerequisites: xLSTM and the mLSTM block
As previously discussed, xLSTM [16] proposes two new building blocks, sLSTM and mLSTM. In line with Vision-LSTM [17], the parallelizable mLSTM module is the fundamental building block of our approach, and thus is the focus of our discussion. The mLSTM block addresses the issues with the original LSTM cell by utilizing a matrix memory cell $C\in\mathbb{R}^{d\times d}$. mLSTM stores a key-value vector pair $k_{t},v_{t}\in\mathbb{R}^{d}$ at timestep $t$, and the relevant value vector is retrieved later using a query $q_{t+\tau}\in\mathbb{R}^{d}$. Equations 1-9 describe the forward pass for mLSTM module, where $C_{t}$ is the matrix memory cell, $n_{t}$ and $h_{t}$ represent the normalizer state and the hidden state, and $i_{t},f_{t}$ and $o_{t}$ represent the input, forget and the output gates, respectively. Weights $W_{q},W_{k}$ and $W_{v}$ are learnable projection matrices for vectors query $q$, key $k$ and value $v$, respectively. Since there is no memory mixing in mLSTMs, multiple memory cells and multiple heads are equivalent, and the forward pass can be parallelized. Further, mLSTM uses exponential gating to mitigate the inability of LSTMs to revise storage decisions when a more similar input is found, which impacts retrieval performance.
@@ 2.1 Prerequisites: xLSTM and the mLSTM block
2.2 AxLSTM: Modelling masked patches with xLSTMs
Creating and masking patches: Given an input spectrogram $\mathbf{x}\in\mathbb{R}^{\mathtt{T}\times\mathtt{F}}$, the axes denoting time and frequency, respectively, we compute non-overlapping patches of shape $t\times f$, yielding $\mathbf{x}_{p}\in\mathbb{R}^{N\times(t\cdot f)}$ patches. Similar to [20, 7], fixed sinusoidal 2d positional embeddings are added to the patches after projecting them to a $\mathbb{R}^{N\times d_{m}}$ dimensional space. To remain in line with previous work [20, 6] and to facilitate fine-tuning of the obtained models, a representative class token is then added to the beginning of the sequence. We then proceed to randomly mask 50% of the input patches using an unstructured masking strategy and replace these masked patches with a learnable mask token. Thus, input to the encoder is:
@@ 2.2 AxLSTM: Modelling masked patches with xLSTMs
Encoding and Reconstruction: These partially masked patches are now fed to the encoder, yielding encoded representations $\mathbf{z}=\mathtt{enc}(\mathbf{x}^{\prime}),\mathbf{z}\in\mathbb{R}^{(N+1)\times d_{m}}$. The encoder is a stack of mLSTM blocks as shown in Fig. 1, where each block expands the $d_{m}$ dimensional input by an expansion factor $E_{f}$ (usually $2$ [16, 17])) before projecting it back to $d_{m}$. The resulting blocks have fewer parameters than a comparable transformer block. To facilitate bidirectional modeling of the input, a flip operation that reverses the order of the input sequence is enabled in even numbered mLSTM blocks, similar to [17]. After encoding, a single hidden layer MLP is used to reconstruct patches from encoded representation $\mathbf{z}$:
@@
where $\mathtt{Linear}_{d}$ is a parameterized linear projection to dimensions $d$, and $\sigma$ denotes the GELU non-linear activation function. Finally, the mean square error (MSE) between the original input spectrogram and the reconstructions, computed only for the masked patches, is used for pretraining. Further details can be found in Section 3.2.
Table 1: Overview of downstream tasks
ID Name Description #Classes #Hours
BO Beijing Opera percussion instrument classification 4 0.3
CD Crema-D emotion recognition 6 10
E50 ESC-50 environmental sound classification 50 2.77
LC LibriCount counting speakers (classification) 10 8
Mri-S Mridangam Stroke classifying Mridangam strokes 10 1.57
Mri-T Mridangam Tonic classifying Mridangam tonics 6 1.57
NS-5h NSynth Pitch 5h pitch classification 88 5.5
SC-5h SpeechCommands 5h keyword spotting 12 6.5
F50K FSD50K multilabel audio tagging 200 100
VL VoxLingua107 Top10 spoken language identification 10 5
SSAST as a Baseline: SSAST [6] is the most directly comparable transformer-based approach to AxLSTMs. As the first patch-based self-supervised learning framework for audio spectrogram transformers (ASTs), it is a well-established method in audio representation learning. While newer transformer-based approaches such as masked autoencoders [21, 7, 22] and self-distilled tokenizers in BEATs [19] introduce additional architectural modifications, SSAST offers a modular framework for evaluating the modeling capabilities of xLSTM v/s the transformer without additional confounding factors. Unlike the original SSAST, which used a combined token prediction and reconstruction based pretraining objective, we simplify the approach by using a reconstruction-only objective.
Table 2: Comparing AxLSTMs with popular self-supervised audio representations. We used pretrained models from cited papers to extract fixed feature vectors and conducted our own downstream experiments. LS, AS, VP, LL stand for LibriSpeech, AudioSet, VoxPopuli and LibriLight datasets, respectively. Original SSAST [6] was trained on AS+LS, whereas we pretrained the directly comparable underlined SSAST baselines (SSAST-T,S,B). *includes decoder parameters
Music & Pitch Speech-based tasks Audio
Model Data #M Params BO Mri-S Mri-T NS-5h CD LC SC-5h VL E50 F50K $s(m)$
Supervised Baselines
HEAR-Naive [23] - - 52.6±2.4 38.0±1.3 36.4±1.9 18.6±4.4 30.9±0.8 33.5±1.1 8.5±0.4 11.2±0.5 5.8±0.2 7.1±0.2 5.1±0.7
PaSST-Base [24] AS $86$ 94.9±0.5 96.5±0.1 87.6±0.6 23.3±0.9 61.0±0.3 60.1±0.2 66.6±1.4 25.5±0.8 94.8±0.3 64.2±0.1 74.4±0.4
@@
W2V2-large [4] VP $315.4$ 93.1±0.7 93.9±0.1 77.4±0.2 42.0±1.0 66.9±0.4 62.4±0.3 87.6±0.5 53.6±1.0 60.1±0.5 34.2±0.1 74.9±0.4
WavLM-large [8] Mix $315.4$ 96.4±0.5 96.8±0.1 89.5±0.1 53.7±0.5 57.2±0.2 61.1±0.3 46.2±0.8 23.7±0.9 47.9±0.4 29.0±0.1 64.8±0.2
HuBERT-large [5] LL $315.4$ 94.1±0.7 95.3±0.1 83.5±0.3 19.3±0.8 70.7±0.1 59.9±0.2 83.2±0.7 66.1±0.9 60.3±0.4 31.5±0.1 74.3±0.3
BEATs-iter3 [19] AS $90.0$ 94.0±0.8 94.7±0.1 95.8±0.1 69.4±0.8 67.3±0.2 68.0±0.2 85.2±0.3 38.5±1.0 83.7±0.3 53.6±0.2 86.7±0.3
AudioMAE [7] AS $86.0$ 93.7±0.6 89.2±0.2 86.6±0.2 64.5±0.8 68.2±0.2 42.2±0.2 28.6±1.5 29.7±1.0 60.6±0.4 37.9±0.1 63.6±0.3
MSM-MAE-208 [21] AS ${92.7}^{*}$ 95.7±0.7 97.3±0.1 97.9±0.1 69.1±0.5 68.7±0.2 63.8±0.5 85.7±0.3 40.3±0.6 78.4±0.6 49.5±0.1 86.0±0.2
MWMAE-Tiny [22] AS ${12.6}^{*}$ 93.3±1.0 97.1±0.1 97.6±0.1 68.1±0.4 64.4±0.2 65.5±0.3 77.0±0.6 28.6±1.1 71.9±0.5 43.4±0.1 80.0±0.3
MWMAE-Base [22] AS ${92.5}^{*}$ 96.0±0.5 97.4±0.1 97.9±0.1 69.3±0.6 73.1±0.3 68.8±0.2 90.9±0.2 44.2±0.9 81.2±0.4 51.2±0.2 90.3±0.2
SSAM-Tiny [14] AS $4.8$ 93.7±0.8 97.1±0.1 94.9±0.1 62.0±0.7 61.8±0.3 59.2±0.4 74.8±0.4 27.8±1.0 70.6±0.2 41.3±0.2 75.6±0.2
SSAM-Small [14] AS $17.9$ 94.0±0.7 97.5±0.1 96.7±0.1 66.3±0.8 67.5±0.2 60.5±0.3 83.7±0.3 39.6±0.7 78.7±0.6 48.5±0.1 83.5±0.3
SSAM-Base [14] AS $69.3$ 93.2±1.1 97.7±0.1 96.9±0.1 70.5±0.5 70.3±0.2 63.5±0.2 87.9±0.3 50.4±0.7 81.0±0.3 52.2±0.1 88.7±0.3
@@
SSAST [6] Mix $89.0$ 93.4±0.9 96.7±0.1 96.3±0.1 66.8±0.7 56.5±0.2 60.7±0.3 53.5±1.3 28.5±0.9 68.4±0.4 38.2±0.1 72.5±0.2
SSAST-T AS $5.4$ 90.4±0.7 95.7±0.1 94.3±0.1 61.2±0.5 46.9±0.2 42.7±0.2 50.6±1.6 13.8±1.0 42.4±0.6 24.6±0.1 55.6±0.2
SSAST-S AS $21.5$ 93.2±0.5 96.2±0.1 95.0±0.1 63.8±0.4 51.6±0.2 50.0±0.3 58.3±1.2 15.6±0.7 50.1±0.6 31.6±0.1 63.0±0.2
SSAST-B AS $85.7$ 93.1±0.7 96.6±0.1 96.2±0.2 64.6±0.8 56.0±0.4 52.9±0.3 66.1±1.0 19.2±0.9 59.6±0.7 37.5±0.1 68.7±0.3
@@
AxLSTM-T AS $4.3$ 93.9±0.7 96.8±0.1 94.9±0.1 62.5±0.6 57.5±0.2 55.1±0.5 69.0±1.6 24.1±0.6 61.3±0.6 37.5±0.2 70.7±0.2
AxLSTM-S AS $16.7$ 92.9±1.0 97.4±0.1 96.6±0.1 66.6±0.4 65.0±0.2 60.3±0.3 80.5±0.4 36.5±0.7 75.5±0.4 46.5±0.1 81.1±0.3
AxLSTM-B AS $65.6$ 93.6±0.9 97.5±0.1 97.5±0.1 71.4±0.8 68.7±0.2 63.2±0.3 85.1±0.2 43.5±0.6 79.2±0.6 51.0±0.1 86.6±0.2
3 Experiments
3.1 Datasets
Pretraining: For pretraining, we use the AudioSet dataset (AS) [18] which has roughly 2 million 10-second long weakly labeled YouTube clips and over 5000 hours of audio data spanning 527 classes.
Downstream Evaluation: Following recent work [22, 14], we use the following subset of tasks proposed as part of the HEAR benchmark [23] for downstream evaluation: Beijing Opera [25, 23], Crema-D [26], ESC-50 [27], LibriCount [28], Mridangam Stroke and Tonic [29], NSynth Pitch 5h [23, 30], Speech Commands 5h [23, 31], FSD50K [32] and VoxLingua107 [33]. Table 1 provides a brief overview of these tasks.
3.2 Implementation details